OpenStudy (clara1223):

find f'(x) of f(x)=sin(2x)(sqrt(1+cos(5x)))

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

product plus chain rule \[(fg)'=f'g+g'f\] with \[f(x)=\sin(x), f'(x)=\cos(x), g(x)=\sqrt{1+\cos(5x)}\] \[g'(x)=\frac{-5\sin(5x)}{2\sqrt{1+\cos(5x)}}\]

OpenStudy (clara1223):

what happened to the 2x inside sin(x)?

OpenStudy (misty1212):

oops \[f(x)=\sin(2x), f'(x)=2\cos(2x)\] my bad

OpenStudy (clara1223):

How would I apply the chain rule to that? The sqrt makes it pretty messy.

OpenStudy (clara1223):

@misty1212 ?

OpenStudy (misty1212):

\[\left(\sqrt{f}\right)'=\frac{f'}{2\sqrt{f}}\]

OpenStudy (misty1212):

which should explain \[g'(x)=\frac{-5\sin(5x)}{2\sqrt{1+\cos(5x)}}\]

OpenStudy (clara1223):

That makes sense. I'm just not sure what to do from there. From the product rule I got 2cos(2x)(sqrt(1+cos(5x)))+sin(2x)((-5sin(5x))/2(sqrt(1+cos(5x))))

OpenStudy (misty1212):

leave it

OpenStudy (clara1223):

that's my final answer?

OpenStudy (misty1212):

it is a silly made up question anyway what else can you do?

OpenStudy (misty1212):

you sure as hell don't want to actually add them although you could if you had like an extra half hour to waste

OpenStudy (clara1223):

Haha definitely not. Still got 6 more questions on this homework sheet due tomorrow.

OpenStudy (misty1212):

best get busy

OpenStudy (misty1212):

lol not that "get busy"